# Jet-shaped filamentary ejecta in common envelope evolution

### v1

@@ 4 Bubbles and filaments in the ejecta
This section describes the evolution when we launch jets from the position of the NS. The simulations do not include the gravity of the NS and reveal the sole roles of the jets. We pre-set the orbit of the NS to spiral-in from $a_{\rm i}=850\,R_{\odot}$ to $a_{\rm SR}=300\,R_{\odot}$ in a timespan of 3 years. It then performs a circular orbit at radius $a_{\rm SR}$. The radial velocity during the rapid in-spiral (plunge-in) phase is constant, while the tangential velocity is the Keplerian one at each radius (details are in Hillel, Schreier, & Soker 2023). The power of the jets is a small fraction of the power that a Bondi-Hoyle-Lyttleton accretion releases; the small factor results from the negative jet feedback mechanism; the power of the jets increases from $\dot{E}_{2j}(a_{\rm i})=1.6\times 10^{41}{~{}\rm erg}{~{}\rm s}^{-1}$ at the outer orbital separation to $\dot{E}_{2j}(a_{\rm SR})=1.4\times 10^{42}{~{}\rm erg}{~{}\rm s}^{-1}$ at the final orbit (details are in Hillel, Schreier, & Soker 2023).
@@
Figure 7 compares the density maps in two planes at three times of three simulations with jets: the left column is the non-rotating simulation as in Hillel, Schreier, & Soker (2023), the middle and right columns are new simulations for a rotating envelope at $t=0$ with values of $\alpha=0.25$ and $\alpha=0.5$, respectively, where $\alpha$ is defined in equation (1). The spiral structure in the planes parallel to the equatorial plane (upper three rows) becomes more prominent with increasing initial envelope rotation (moving from left to right). Other than this, the general qualitative appearance of the density maps is similar in the three simulations. Therefore, we will not repeat the study of the ejecta morphology, the shedding of pairs of vortices in an expanding spiral pattern as the NS spirals-in, and the entropy profiles that we explored for the non-rotating simulation in Hillel, Schreier, & Soker (2023).
@@
Figure 7: Density maps of the three simulations with $\alpha=0$, $0.25$ and $0.5$, left to right, at $t=3.8{~{}\rm yr}$. The first three rows are for different planes parallel to the equatorial plane; the insets show the distance of each plane from the equatorial plane in units of $10^{12}{~{}\rm cm}$, namely, $z=6\times 10^{12}{~{}\rm cm}$, $21\times 10^{12}{~{}\rm cm}$, and $100\times 10^{12}{~{}\rm cm}$. The lower row shows the density maps in the plane $x=0$. The initial (at $t=0$) NS location is marked with a black dot on the upper left panel. These density maps emphasize stronger spiral structures for faster rotation and the formation of medium-size structures, i.e., filaments and bubbles.
@@
Figure 1 shows that the large energy that the jets deposit to the envelope causes its expansion to the degree that the shells we follow do not perform oscillations (solid lines in the two panels of the figure).
@@
$\tau_{\rm RT}\simeq\frac{\rho}{\sqrt{-\overrightarrow{\nabla}P\cdot\overrightarrow{\nabla}\rho}}.$ (2)
To present both stable and unstable zones in the maps to follow, we refer to the growth rate of the instability
$f_{\rm st}\equiv\frac{1}{\rho}{\sqrt{\left|\overrightarrow{\nabla}P\cdot\overrightarrow{\nabla}\rho\right|}}~{}\text{sgn}(\overrightarrow{\nabla}P\cdot\overrightarrow{\nabla}\rho).$ (3)
In these maps, negative values show the Rayleigh-Taylor unstable zones, and the magnitude of $f_{\rm st}$ is the growth rate of the instability, i.e., $-1/f_{\rm st}$ is the growth time. In contrast, positive values of $f_{\rm st}$ are stable zones, and the value is about the Brunt–Väisälä frequency.
@@ 5 Summary
We do not have the computer resources to perform the simulations with jets in the high-resolution grid. The comparison of the regular and high-resolution simulations without jets (Figure 2 - 6) teaches us that the regular resolution is adequate to perform the simulations with jets, at least for our goals.
Figures 7 and 8 present the results of the simulations with the jets that the in-spiraling NS launches. Figure 7 compares the density of the bounds and ejected gas in the non-rotating simulation with two rotating envelope simulations. The jets’ power is the same in the three simulations. The most obvious difference is that the faster the rotation is, the more pronounced the spiral structure in planes parallel to the orbital plane is. The general conclusion, however, is that there are no significant differences between the simulations with different envelope rotations for the jets’ power we employed. We, therefore, do not present more properties of the ejected gas that we presented for the non-rotating case in Hillel, Schreier, & Soker (2023).
A new analysis of this study is that of Rayleigh-Taylor instabilities. Figure 8 presents the unstable zones in two planes and two times for the $\alpha=0.5$ simulation (equation 1 defines $\alpha$). The simulation with slower rotation and the one without envelope rotation are also prone to Rayleigh-Taylor instabilities, as the instability results from the energy deposition by jets. These instabilities account for the filaments and bubbles of the spiral structure and the ejected gas that figure 7 presents. Clumpy ejecta (with filaments, bubbles, and clumps) influences the light curve from such events and should be the subject of future studies.
In a broader context, jets seem to be the most prominent observable of CEE, at least in planetary nebulae with central binary stars (Soker, 2025). Studying the jets’ roles in CEE should be a prime focus of CEE research. Our new study adds to the exploration of the role of jets in revealing some properties of Rayleigh-Taylor instabilities that jets induce.
@@ Acknowledgments
We thank Aldana Grichener for her valuable comments. The Amnon Pazy Research Foundation supported this research.

### v2

@@ 4 Bubbles and filaments in the ejecta
This section describes the evolution when we launch jets from the position of the NS. The simulations do not explicitly include the gravity of the NS and reveal the sole roles of the jets. Namely, the main role of the NS gravity is in accreting mass and launching jets. The NS gravitational force on the envelope is not taken into account. We explore the role of jets in this CEE process. We pre-set the orbit of the NS to spiral-in from $a_{\rm i}=850\,R_{\odot}$ to $a_{\rm SR}=300\,R_{\odot}$ in a timespan of 3 years. It then performs a circular orbit at radius $a_{\rm SR}$. The reason we stop the spiraling-in of the NS at $a_{\rm SR}=300\,R_{\odot}$ is numerical. To conserve computational resources, we do not simulate the evolution within the inner $20\%$ of the envelope, specifically in the inert core, which has a radius of $R_{\rm inert}=176R_{\odot}$. We cannot deposit the energy of the jets too close to this radius and, therefore, do not allow the NS to spiral in below $a_{\rm SR}=300R_{\odot}$. The radial velocity during the rapid in-spiral (plunge-in) phase is constant, while the tangential velocity is the Keplerian one at each radius (details are in Hillel, Schreier, & Soker 2023). The power of the jets is a small fraction of the power that a Bondi-Hoyle-Lyttleton accretion releases; the small factor results from the negative jet feedback mechanism; the power of the jets increases from $\dot{E}_{2j}(a_{\rm i})=1.6\times 10^{41}{~{}\rm erg}{~{}\rm s}^{-1}$ at the outer orbital separation to $\dot{E}_{2j}(a_{\rm SR})=1.4\times 10^{42}{~{}\rm erg}{~{}\rm s}^{-1}$ at the final orbit (details are in Hillel, Schreier, & Soker 2023).
@@
Figure 7 compares the density maps in two planes at three times of three simulations with jets: the left column is the non-rotating simulation as in Hillel, Schreier, & Soker (2023), the middle and right columns are new simulations for a rotating envelope at $t=0$ with values of $\alpha=0.25$ and $\alpha=0.5$, respectively, where we defined $\alpha$ in equation (1). The spiral structure in the planes parallel to the equatorial plane (upper three rows) becomes more prominent with increasing initial envelope rotation (moving from left to right). Other than this, the general qualitative appearance of the density maps is similar in the three simulations. Therefore, we will not repeat the study of the ejecta morphology, the shedding of pairs of vortices in an expanding spiral pattern as the NS spirals-in, and the entropy profiles that we explored for the non-rotating simulation in Hillel, Schreier, & Soker (2023).
@@
Figure 7: Density maps of the three simulations with $\alpha=0$, $0.25$ and $0.5$, left to right, at $t=3.8{~{}\rm yr}$. The first three rows are for different planes parallel to the equatorial plane; the insets show the distance of each plane from the equatorial plane in units of $10^{12}{~{}\rm cm}$, namely, $z=6\times 10^{12}{~{}\rm cm}$, $21\times 10^{12}{~{}\rm cm}$, and $100\times 10^{12}{~{}\rm cm}$. The lower row shows the density maps in the plane $x=0$. The upper left panel shows a black dot marking the initial (at $t=0$) NS location. These density maps emphasize stronger spiral structures for faster rotation and the formation of medium-size structures, i.e., filaments and bubbles.
@@
Figure 2 shows that the large energy that the jets deposit into the envelope causes its expansion to the degree that the shells we follow do not perform oscillations (solid lines in the two panels of the figure).
In Figure 8 we compare the density maps of three simulations with three resolutions for $\alpha=0.5$, and at the same time and planes as in Figure 7. The middle column of Figure 8 is the same as the right column of Figure 7; the left column is for lower resolution and the right column is for higher resolution. The lower resolution (left column) does not show a clear spiral pattern. Other than this, all simulations show the formation of filamentary extended envelope and ejecta.
Figure 8: Density maps of three simulations with different resolutions, all with $\alpha=0.5$. The density maps are at $t=3.8{~{}\rm yr}$ and in the same four planes as in Figure 7. The first three rows are for different planes parallel to the equatorial plane; the insets show the distance of each plane from the equatorial plane in units of $10^{12}{~{}\rm cm}$, namely, $z=6\times 10^{12}{~{}\rm cm}$, $21\times 10^{12}{~{}\rm cm}$, and $100\times 10^{12}{~{}\rm cm}$. The lower row shows the density maps in the plane $x=0$. The upper left panel shows a black dot marking the initial (at $t=0$) NS location. The middle column is the same as the right column of Figure 7, where all cells have a size of $\Delta_{\rm cell}=3.90625\times 10^{12}{~{}\rm cm}$. The left column has a lower resolution with a larger grid-cell size by a factor of $1/0.714=1.4$. The right column has a higher resolution with a grid-cell size smaller by a factor of $1/1.25=0.8$; the grid is smaller due to the small cell size.
@@
$\tau_{\rm RT}\approx\frac{\rho}{\sqrt{-\overrightarrow{\nabla}P\cdot\overrightarrow{\nabla}\rho}},$ (2)
where we assume that the wavelength of the perturbation is about the density scale height, $(d\ln\rho/dr)^{-1}$ (e.g., chapter 7 in Priest 1982). To present both stable and unstable zones in the maps to follow, we refer to the growth rate of the instability
$f_{\rm st}\equiv\frac{1}{\rho}{\sqrt{\left|\overrightarrow{\nabla}P\cdot\overrightarrow{\nabla}\rho\right|}}~{}\text{sgn}(\overrightarrow{\nabla}P\cdot\overrightarrow{\nabla}\rho),$ (3)
which is a frequency in stable zones (hence the letter ‘f’ and the subscript ‘st’). In unstable zones $f_{\rm st}<0$, and $-1/f_{\rm st}$ is approximately the growth time of the RTI (equation 2). In stable zones $f_{\rm st}$ is approximately the Brunt–Väisälä frequency (e.g., chapter 4 in Priest 1982), assuming the density gradient is much steeper than the pressure gradient, $|d\ln\rho/dr|\gg|d\ln P/dr|$; the latter holds in our case where the low-density zone are much hotter.
@@ 5 Summary
We do not have the computer resources to perform the simulations with jets in the high-resolution grid. The comparison of the regular and high-resolution simulations without jets (Figure 1, and 3 - 6) teaches us that the regular resolution is adequate to perform the simulations with jets, at least for our goals.
Figures 7 and 9 present the results of the simulations that include the jets that the NS launches inside the RSG envelope. Figure 7 compares the density of the bounds and ejected gas in the non-rotating simulation with two rotating envelope simulations. The jets’ power is the same in the three simulations. The most obvious difference is that the faster the rotation is, the more pronounced the spiral structure in planes parallel to the orbital plane is. The general conclusion, however, is that there are no significant differences between the simulations with different envelope rotations for the jets’ power we employed. We, therefore, do not present more properties of the ejected gas, which can be found in our study of the non-rotating case in Hillel, Schreier, & Soker (2023).
A new analysis of this study is that of Rayleigh-Taylor instabilities. Figure 9 presents the unstable zones in two planes and two times for the $\alpha=0.5$ simulation (equation 1 defines $\alpha$). The simulation with slower rotation and the one without envelope rotation are also prone to Rayleigh-Taylor instabilities, as the instability results from the energy deposition by jets. These instabilities account for the filaments and bubbles of the spiral structure and the ejected gas that figure 7 presents. Clumpy ejecta (with filaments, bubbles, and clumps) influence the light curve from such events and should be the subject of future studies.
In a broader context, jets seem to be the most prominent observable of CEE, at least in planetary nebulae with central binary stars (Soker, 2025). Studying the roles of jets that the secondary star launches in CEE should be a prime focus of CEE research. Our new study adds to the exploration of the role of jets in revealing some properties of Rayleigh-Taylor instabilities that jets induce.
@@ Acknowledgments
We thank Aldana Grichener for her valuable comments. We thank an anonymous referee for detailed comments and suggestions. The Amnon Pazy Research Foundation supported this research.